The length of a rectangle is five times its width. The area of the rectangle is 180 square centimeters. Find the dimensions of the rectangle.

Answer by checkley79(3341)

L=5W
A=LW
180=5W*W
10=5W^2
W^2=180/5
W^2=36
W=SQRT36
W=6 ANS. FOR THE WIDTH.
L=5*6=30 ANS. FOR THE LENGTH.
PROOF:
180=30*6
180=180